A group of 5 people are going to meet weekly at the library for 4 weeks. Each week, two people are selected at random to speak. Each person may speak in multiple weeks, but no pair of people will speak together more than once. The probability that there is a person who will never be asked to speak can be expressed as a/b where a and b are co-prime positive integers. What is the value of a+b?

Grade:11

1 Answers

Shivam Dimri
11 years ago

Well probability!!1

i like it,,

okay we have 5 people here,

total pairs possible  is 5C2 = 10

suppose 1 person will not speak then we can select that person in 5 ways

remaining pairs possible = 4C2 = 6

Now as the pairs will be speaking for 4 weeks we have to ensure the same pair does not speaks!!

so total selections possible without any restrictions = 10C4 

and total favourable selections = 5  *  6C4

so i believe you can find out the probablity = Favourable / total

it comes out to be  = 5/14

sum = 5 + 14 = 19

thats my lucky number

do reply....